Cos(5x)cos(2x) + sin(5x)(sin(2x)


1
Expert's answer
2021-02-24T07:12:28-0500

Given: "cos(5x)cos(2x)+sin(5x)sin(2x)"

Recollect the following:

"cos(A-B)=cosAcosB+sinAsinB"

Using the above, we have

"cos(5x)cos(2x)+sin(5x)sin(2x)=cosAcosB+sinAsinB, A=5x,B=2x"

"=cos(A-B)"

"=cos(5x-2x)"

"=cos(3x)"

Therefore, "cos(5x)cos(2x)+sin(5x)sin(2x)=cos(3x)"